Thomson Trucking has $21 billion in assets, and its tax rate is 35%. Its basic earning power (BEP) ratio is 17%, and its return on assets (ROA) is 3%. What is its times-interest-earned (TIE) ratio? Round your answer to two decimal places.
Answer #1

Please provide rating.. Let me know if you need any clarification ..

Given that-
BEP = 17%
Which mean
EBIT/Total asset = 17%
Also Total asset = 21 Billion
Therefore EBIT = 21billion *17% 3.57 Billion
ROA is 3%
ROA = Net income/Total asset
Net income = Total asset *3% =21Billion*3%          0.63 billion
Earning before tax = Net income/(1-Tax) = 1/(1-35%)          0.97 billion
Interest expenses =EBIT-EBT = =3.57-.97          2.60 billion
Times interest earned ratio = EBIT / Interest expense =3.57/2.6          1.37 times